# user interface, color vision, color blindness, CVD, accessibility

### 18 Jan 2024 v1

@@ Abstract
Color vision deficiency (CVD, color blindness) is the failure or decreased ability to distinguish between colors under normal lighting conditions. There are over 300 million people worldwide with CVD, including approx. 1 in 12 men (8%) and 1 in 250 women (0.5%). CVD can limit a user’s ability to interact with websites and software packages that are otherwise basic commodities. User interface designers have taken various approaches to tackle the issue with some interfaces offering a high contrast mode, and others integrating color-blind awareness into their website design process. The Web Content Accessibility Guidelines (WCAG) outline some best practices for maintaining accessibility that have been adopted and recommended by several governments; however, it is currently uncertain how this impacts perceived user functionality and if this could result in a reduced aesthetic look. In our work, we present a subjective user study to measure the loss of functionality and aesthetics as potentially seen by CVD observers for 20 popular websites and software packages. As recruiting participants with CVD is non-trivial, we developed a simulation-based pipeline instead for our full-reference mean opinion score experiment, which as far as we know is a novelty in the field. Our results show that relative aesthetics and functionality correlate positively and that an operating-system-wide high contrast mode can reduce both aesthetics and functionality for CVD users. Finally, we propose a AAA–A classification of the interfaces we analyzed.
@@ 1 Introduction
Color vision deficiency (CVD, color blindness) is the failure or decreased ability to distinguish between colors under normal illumination. There are over 300 million people with CVD, including approx. 1 in 12 men (8%) and 1 in 250 women (0.5%) Birch (2012); Rigden (1999); Al-Aqtum and Al-Qawasmeh (2001). CVD is an X-linked genetic disorder impacting both eyes with varying degrees of prevalence in different populations Fakorede et al. (2022). It affects an individual’s ability to perform tasks in both personal and professional settings Tagarelli et al. (2004).
@@
Color is an important asset in user interface (UI) design wca (2018), and while the exact impact of color is known to vary between demographics Swasty and Adriyanto (2017), applications can still often rely on established conventions, such as green and red indicating ‘yes‘ and ‘no’ respectively. Objects of the same color satisfy the Gestalt principle of similarity, whereas different colors can help an object stand out or mark figure-ground articulation Todorovic (2008). With the ever-increasing color gamut of novel displays Wen et al. (2021), new domains are opening up in the use of color; however, some of these domains simply cannot be seen by someone with CVD.
@@
UI designers have taken various approaches to tackle the issue of CVD, including the use of a reduced color palette (high contrast mode), integrating CVD awareness into the design process, or relying on automatic post-processing enhancements (daltonizationSimon-Liedtke and Farup (2016). The Web Content Accessibility Guidelines (WCAG) wca (2018) outline some best practices for maintaining accessibility; however, these often only target core functionality. Another consideration for UI is perceived aesthetics. Many professional designers see aesthetics as inversely proportional to functionality Karvonen (2000), while other evidence points towards a positive correlation between functionality and aesthetics Ahmed et al. (2009); Kurosu and Kashimura (1995); Tractinsky (1997). In the domain of CVD, a high contrast theme is rarely a top-priority feature, which could imply that people with CVD have a substantially reduced aesthetic experience. However, there is generally insufficient existing data to understand CVD users’ perceived functionality and aesthetics of UIs.
In this work, we conducted a subjective user study with 19 participants (1449 data points) to measure the perception of 20 popular UIs. Specifically, we measured mean-opinion scores for functionality and aesthetic loss in CVD screenshots with respect to the original reference images. To obtain relative scores between the original UI and the one perceived by CVD observers, we developed a novel, simulation-based pipeline that could be also used by UI professionals. Our main contributions can be summarized as follows:
@@
  • Collection of subjective user data on the loss of functionality and aesthetics when seen through CVD simulation
@@
The rest of the paper is structured as follows: in Section 2, we discuss the relevant background literature, in Section 3 we describe the proposed experimental methodology, before discussing our findings in Sections 4 and 5.
@@ 2 Background
2.1 Types of CVD
Human color vision is widely considered trichromatic due to the three types of retinal cone cells (L, M, and S) that respond differently to wavelengths of light (Fig. 1) Wandell (1995). According to opponent color theory Hurvich and Jameson (1957), higher-order visual functions rely on a differential response of cones, with luminance (L+M) being more prominent than chrominance channels (L-M, L+M-S). CVD can occur as a result of cone cells being absent, not working, or having an abnormal response. Severe CVD occurs when two or three types of cone cells are absent (monochromacy, achromatopsia). Less severe CVD occurs when one type of cone is absent (L: protanopia, M: deuteranopia, S: tritanopia), or all three are present, but one cell has an abnormal response (L: protanomaly, M: deuteranomaly, S: tritanomaly) Huang et al. (2008). Red-green CVD, specifically protanomaly and deuteranomaly, are the most frequent Marey et al. (2015), which can be intuitively explained by the spectral similarity of the M and L cones. This affects approx. 8% of the male and 0.1–0.3% of the female population Birch (2012).
Figure 1: Normalised spectral sensitivity of the three types of retinal cone cells as a function of light wavelength. Note the similarity of the $M$ and $L$ cones, where minor individual differences can result in an unresolvable overlap, causing observers to have a reduced ability to differentiate between lights in the red-green region. Plotted based on Stockman and Sharpe Stockman and Sharpe (2000).
@@ 2.2 WCAG
Accessibility is the concept of making UIs equally usable by all types of users, enabling all user interactions without barriers. In our work, we consider functionality to be synonymous with usability, therefore accessibility also aims to maintain functionality. In the past few decades, accessibility has become a legal requirement in several countries Easton (2012), with most government agencies recommending the adoption of the World Wide Web Consortium’s (W3C) Web Content Accessibility Guidelines (WCAG) Easton (2012); wca (2018). In fact, WCAG is consulted by website and UI developers every day to enforce best practices to maintain a level playing field for all users, regardless of age, or physical or mental disabilities. WCAG has also been utilized in academic research to assess website accessibility Al-Khalifa (2010); Kuzma (2010); Paul (2023)
WCAG 2.1 wca (2018), the latest framework as of now, specifies a set of success criteria, with an associated AAAA system, A being the least and AAA being the most restrictive. The minimum accessibility compliance is often set at AA. Easton (2012).
CVD support is addressed across several success criteria, including:
@@
It remains uncertain to what extent current UIs implement these recommendations. Also, while the use of contrast ratio thresholds gives designers an intuitive scheme, contrast is defined as a color difference perceived by trichromatic (non-CVD) observers, which is not trivially convertible to contrast for CVD users. Furthermore, WCAG contrast is typically calculated for background vs. foreground, which can fail to capture some impacts of CVD, such as reduced differences across UI components (Figure 2). We accept the importance and the positive impact of WCAG, but for CVD, we argue that there is a need for a subjective user study to measure accessibility.
Figure 2: Illustration of a popular quiz platform’s UI processed with simulated tritanomaly Machado et al. (2009). While text vs. background contrast remains mostly unchanged, $\Delta E_{2000}$ color difference Luo et al. (2001) between buttons 1 and 2 is reduced from $25.1$ to $14.1$. Similarly, $\Delta E_{2000}$ between buttons 3 and 4 is reduced from $43.0$ to $18.3$. The colors remain distinguishable, but the difference between them is less noticeable. WCAG analysis would flag an insufficient text vs. background contrast on button 3 (<3), but current guidelines do not address the reduced color difference between buttons.
@@ 2.3 Reducing the impact of CVD
To reduce the impact of CVD, WCAG encourages UI designers to convey information via alternate visual channels. This also agrees with recent research on the use of patterns Sajadi et al. (2012), and textures Wu et al. (2019) to improve accessibility.
Another widely adopted approach limits the color palette with an increased contrast between key colors. E.g. Windows users can select from multiple operating-system wide high contrast themes that accommodate different types of vision deficiencies. While almost all applications and websites respect the high contrast theme of the operating system, the resulting UI often looks substantially different, with seemingly reduced information for a non-CVD observer (see Figure 3). In our study, we investigate whether this strategy is viable by measuring functionality and aesthetics scores with the operating system’s high contrast mode enabled and disabled.
An alternative technique is daltonization, an umbrella term for image-based recoloring techniques that enhance displayed images for CVD observers Ribeiro and Gomes (2019); Machado and Oliveira (2010). Such algorithms aim to alter colors that are isochromatic to CVD observers but are perceived differently by a non-CVD observer. Daltonization could substantially improve CVD accessibility; however, current algorithms are still limited by e.g. run-time cost, lack of temporal coherence, or the ability to accommodate different types and levels of CVD Ebelin et al. (2023).
@@
2.4 Aesthetics
While accessibility and overall functionality are key features of a good UI, designers, and human-computer interaction (HCI) researchers are increasingly discovering the impact of attractiveness and beauty. An attractive UI can help make a connection between a user and the application Türkyilmaz et al. (2015). Interdisciplinary research has investigated the relevance of aesthetics, a classical philosophical field focusing on the “perception of the beautiful in nature and art” Oxford English Dictionary (2023). A particular focus has been the difference between classical and expressive aesthetics in HCI, with classical aesthetics referring to the traditional notions of an orderly and clear design (e.g. sizing), and expressive aesthetics referring to creativity and originality in design Hartmann et al. (2008). Mahlke and Thüring Mahlke and Thüring (2007) argue that classical aesthetics is perceived more evenly, whereas expressive aesthetics varies based on framing effects; as such, in this paper, we focus on classical aesthetics. For a more complete review of the literature, refer to Ahmed et al. Ahmed et al. (2009).
The relationship between UI aesthetics and functionality has been investigated by a number of authors. Some argue that the two goals often conflict Karvonen (2000), whereas Kurosu and Kashimura Kurosu and Kashimura (1995) found a positive correlation between the beautifulness of ATM interfaces and their usability. This positive correlation has been since verified by other publications, but the correlation coefficient has been shown to be culturally dependent Tractinsky (1997). To our knowledge, no previous publication has measured aesthetics for CVD users. We hypothesize that that functionality and classical aesthetics correlate, and classical aesthetics could be negatively impacted by automatic invasive techniques such as the high contrast mode.
@@ 3 Materials and Methods
To explore the impact of CVDs, we conducted a subjective user experiment. Specifically, we investigated the following two aspects:
@@
  • Aesthetics: according to classical aesthetics, is the clarity and orderliness of the UI retained when viewed by a CVD observer?
In our experiment design, we consider the three most common types of CVDs: protanomaly, deuteranomaly, and tritanomaly. As recruiting participants, especially with tritanomaly, is challenging, and a large sample would be required to establish the relative functionality and aesthetic scores by CVD vs. non-CVD users, we propose a full-reference simulation-based experiment, where participants without CVD can make a comparative judgment between the stimulus as seen by CVD and non-CVD users. Such a pipeline has the added benefit that it is suitable to be deployed in a commercial environment.
@@
In the following subsections, we describe the stimuli, the setup, the task, and the participant sample set.
3.1 Stimuli
To understand if there is a loss of functionality and/or aesthetics, we presented users with the original reference image as well as three distorted images that were identical in terms of content, showing what a CVD person might see (Figure 4). The reference image was always presented in the top-left; participants were made aware of this during the briefing, and there was a “Reference” label to remind them during the experiment. The distorted images were generated using CVD simulation from Machado et al. Machado et al. (2009) as implemented in Smith (2018) assuming an sRGB display. Three CVDs were considered: protanomaly (top-right), deuteranomaly (bottom-left), and tritanomaly (bottom-right). For the experiment, we set the simulation level conservatively to 100%.
For content, we sampled 20 software across different industries (social, business, entertainment, music, shopping, food and beverage, travel, education, productivity, and SaaS). Within each sector, we picked the application or website based on general popularity in the UK, according to informal surveys. This does not give full coverage of all available design techniques and CVD mitigations; however, it can serve as a representative sample of UIs the average user is likely to encounter. The list of UIs: Amazon, Booking.com, Brainly, Dropbox, Duolingo, Facebook, Google Calendar, Google Maps, Indeed, Instagram, LinkedIn, Photomath, Pocket, Slack, Spotify, Teams, Tiktok, Todoist, Uber Eats, and Zoom.
Participants viewed static screenshots. Specifically, in each software, we chose a common user journey that was relevant to the use of that software and recorded 3 screenshots. Some of these screenshots were resized with a maintained aspect ratio, and users could examine the original full-resolution content by hovering with the mouse.
Figure 4: Visualisation of the experiment setup. Top-left image is the reference screenshot, with three types of CVD (protanomaly: top-right, deuteranomaly: bottom-left, tritanomaly: bottom-right) simulated. The user controls the aesthetics and functionality scores at the bottom of the screen.
To understand the functionality and aesthetics of high contrast modes, we presented each application twice (1): distorted images calculated from the reference image, (2): distorted images calculated from high contrast screenshots (but the reference left unaltered). The high contrast images were captured manually, although in the future this could also be automated. We used the Windows Surface High Contrast #1 and #2 themes for protanomaly and deuteranomaly simulation respectively. The only exception was Slack, which has native themes for CVD.
In total, there were 120 stimuli (20 UIs $\times$ 3 screenshots $\times$ 2 modes of standard or high contrast mode).
3.2 Setup
The experiment was presented on a website. Participants used their own devices, which introduces some variability in terms of screen size, viewing conditions, and color gamut. We postulate that screen size and viewing condition differences are representative of the real world; however, gamut differences could increase measurement noise.
3.3 Task
Once presented with the reference and the three distorted screenshots, participants were asked to rate the following (quotes from the briefing form):
  • functionality on a scale of 1–5: “could you use the distorted user interface the same way as the original reference”?
  • aesthetics as a binary lost/kept: “do the distorted images maintain the aesthetics of the original reference? Aesthetics in this context will be defined as maintaining the same clarity and orderliness as the reference” (c.f. classical aesthetics). Our preliminary results indicated that participants found the binary scale easier in this instance.
For every new stimulus, both response sliders were initialized to a random location. Participants had unlimited time to adjust the sliders before pressing the submit button.
3.4 Participants
19 people took part in the experiment (from the UK, age 17–50, 12 male and 7 female). All participants had full-color vision (verified with Ishihara test Clark (1924)). Participation was voluntary without financial incentive, following informed consent. Participants were encouraged to complete all 120 trials; however, this was not a requirement. 7 participants completed all 120 trials, with an average of 79 trials across all participants. As the order of the trials was randomized for each person, this still resulted in a balanced dataset across applications.
4 Results
In this section, we summarise the quantitative results of the subjective user experiment, with a more elaborate discussion in Section 5.
4.1 Functionality
Functionality scores of each UI, averaged over participants and screenshots, is summarised in Figure 5. UIs are ordered from least functionality lost to most functionality lost. Error bars indicate 95% confidence intervals. Different UIs have been rated differently, with a statistically significant difference across the board. However, average functionality scores for all apps are above 3 (on a scale of 5), with the highest scores (Photomath, Zoom, GOogle Calendar) slightly above 4. The standard UI and high contrast mode have been rated very similarly by the participants. For most UIs, there is no significant difference; however, wherever there is ($p<0.05$, indicated with an *), it seems to be in favor of the standard UI.
4.2 Aesthetics
The aesthetics score of each UI, averaged over participants and screenshots, is summarised in Figure 6. There is a more noticeable difference in the aesthetics scores across the board, with average scores ranging from $0.22$ to $0.61$. Furthermore, there is a marked difference between the standard UI, and the high contrast stimuli, with a general preference for the standard UI. The difference is significant ($p<0.05$) for 12 out of the 19 UIs tested.
Figure 5: Average functionality score for each UI. Colors show all responses, standard (non-high-contrast), and high contrast responses. Error bars indicate 95% confidence intervals, asterisks (*) indicate a significant difference between standard and high contrast responses for a given app.
Figure 6: Average aesthetics score for each UI. Colors show all responses, standard (non-high-contrast), and high contrast responses. Error bars indicate 95% confidence intervals, asterisks (*) indicate a significant difference between standard and high contrast responses for a given app.
4.3 Functionality vs. aesthetics
When separating the dataset into standard (non-high-contrast) and high contrast modes, results show a strong correlation (Pearson r=0.74) between the aesthetics score and the functionality score (both averaged over screenshots and participants). Figure 7 contains an overview of the results, with more detail available in Figure 10 in Appendix A.
4.4 High contrast mode
Results indicate that Windows high contrast mode reduces both UI functionality and aesthetics significantly ($p\ll 0.05$).
Figure 7: Average functionality scores of simulated CVD screenshots vs. average aesthetics scores. Each point indicates an app in either high-constrast mode (orange crosses) or in standard (non-high-contrast) mode (blue circles). The dataset shows a strong positive correlation (Pearson r=0.74)
5 Discussion
The simulation-based experiment has proved to be an effective way to gather information on the impact of CVD in the context of user interfaces. However, it is worth noting that functionality and aesthetic scores are relative to the reference images, and give no basis to compare the absolute functionality or aesthetics of the applications.
5.1 Functionality
Results indicate that people with CVD would find it somewhat harder to use the tested applications. This could potentially mean increased reaction time or minor difficulties with interaction; however, the overall scores were favorable (mean=3.88). The positive impact of WCAG seems observable, as most UIs convey the same information using multiple visual clues (c.f. WCAG 1.4.1), and there is good foreground-background contrast. Analyzing the lower functionality score screenshots reveals a few common pitfalls. Firstly, when color is used to draw the user’s attention, the impact might be more subtle for a CVD user (see Figure 8). Some UIs seem to use color to differentiate between groups of similar items (e.g. icons), which can make search tasks harder. The final issue we found seems domain-specific: while many UIs use photos/videos to enrich their design, social media applications rely on these almost exclusively. Such content is naturally more information-rich, and often user-created, which makes it harder to observe accessibility guidelines. E.g. colored text bubbles over a video can be harder to notice by a CVD observer. In the extreme, Instagram has received the lowest functionality score in the experiment, which we postulate is because many of Instagram’s popular image filters appear mostly identical on the CVD screenshots.
Figure 8: Color-coded error field. The UI meets WCAG guidelines with the use of the warning icon; however, the shade of red when seen by a CVD user (deuteranomaly) stands out less, which can make the user interaction slower in a large form.
5.2 High contrast mode
Windows high contrast mode does not seem suitable for maintaining the aesthetics or the functionality of a UI when observed by CVD users. The former is expected, as the high contrast mode alters the UI with e.g. borders and color contrast that are meant to aid accessibility purely. However, the latter is surprising, as one would expect such accessibility tools to increase functionality rather than decrease it. We speculate that there could be two reasons for the low functionality scores: 1) many interfaces adapt to high contrast themes poorly, and key visual information can be lost (e.g. see Figure 9), and 2) if there is a linear relationship between aesthetics and functionality, the poor aesthetics scores can have a negative cognitive impact on the perceived functionality as well.
In our sample, the only application, where enabling a high contrast theme had a noticeably positive impact on aesthetics was Slack. However, Slack is also the only application we tested with its own built-in CVD themes (rather than Windows high contrast). This might imply that applications with their own CVD mode could outperform their competitors in terms of aesthetics (when used by someone with CVD). Future research could verify this with a more balanced design of UIs with and without their custom CVD themes.
Figure 9: Reduced functionality in high contrast mode. In this instance, turning the calendar icon black would have helped to maintain functionality.
5.3 Aesthetics and functionality
Results show a strong linear correlation between functionality and aesthetics for the CVD-simulated screenshots when separating the data set into standard and high contrast modes. While our scores are relative, they seem to agree with the hypothesis that there could be a linear relationship between classical aesthetics and functionality. As such, a UI that is tested for CVD functionality for can be expected to maintain aesthetics as well.
The only strong outlier in the functionality vs. aesthetics plot was Instagram, with a reasonably high aesthetics score, and a surprisingly low functionality score. As discussed, this could be attributed to the inherently challenging task of maintaining functionality in an application, where applying recoloring filters is a core feature.
5.4 Classification framework
Having measured the loss of functionality and aesthetics in a subjective study for a number of UIs, and inspected the relevant screenshots, we propose a simple 3-level classification model to categorize UIs:
  • AAA functionality and aesthetics as seen by CVD users are almost distinguishable from that of a non-CVD user. In our results, these are the applications with high functionality scores (>4) and high aesthetics scores (>0.75)
  • AA aesthetics and functionality are mostly kept for CVD users (functionality > 3.8, aesthetics >0.5)
  • A aesthetics might be lost, and functionality is reduced, but core functionality is still accessible (e.g. through visual means other than color). Meets WCAG AA accessibility guidelines.
Our classification of the UIs from the experiment is shown in Table 1.
6 Limitations
Analyzing all available UIs in a subjective user experiment is infeasible, and the biggest limitation of our work is the sample set and bias. Specifically, many popular UIs we analyzed follow similar designs. Evaluating accessibility and aesthetics in a different field (e.g. video games) would require a new set of experiments. Our use of screenshots means that the results are only representative of a subsection of the application. Future studies could consider an interactive setup. Our sample was based in the UK, and aesthetics scores are known to be culturally dependent. Finally, our simulation-based approach can be a powerful way to estimate how a CVD user might perceive a UI. However; 1) simulations have known limitations, such as taking place in trichromatic space rather than spectral space – this could alter the accuracy of the simulation 2) when judging CVD-simulated images, observers could be biased to object to artifacts such as discolorations.
Table 1: Tabular results of our user interface classification. Software ranked from Level A (lowest) to AAA (highest), where the ordering considers both aesthetics and functionality. Tiles on the same level represent a similar overall score.
7 Conclusions
CVD affects over 300 million people, but our understanding of how it impacts user interface (UI) interaction is still limited. In this paper, we measured if a CVD observer would lose out on any of the functionality or the aesthetics of 20 popular applications/websites. The stimuli dataset and the experiment responses are available for other researchers. Our results indicate that most UIs maintain a reasonably good level of functionality and aesthetics. UIs that rely on color to distinguish icons or indicate errors might be harder to use for CVD users. Furthermore, we found a linear relationship between functionality and aesthetics scores, which can simplify UI designers’ tasks when planning for CVD support.
Unfortunately, we found the Windows high contrast mode to be detrimental to the functionality and aesthetics of applications for CVD users. Applications might benefit from implementing their own custom CVD themes, but the significance of this should be tested in future studies.
[…]
This research received no external funding.
[…]
The study was conducted in accordance with the Declaration of Helsinki, and approved by the Head of Research of The Perse School Cambridge (18th November 2023).
[…]
Informed consent was obtained from all subjects involved in the study.
[…]
The data presented in this study are openly available on github: https://github.com/gdenes355/cvd_ui_dataset
Acknowledgements.
[…]
Appendix A
Figure 10: Average relative functionality scores of simulated CVD screenshots plotted over average aesthetics scores. Each point indicates an app in either high contrast mode (orange crosses) or in standard (non-high-contrast) mode (blue circles).
[…]

### 24 Feb 2024 v2

@@ Abstract
Over 300 million people who live with color vision deficiency (CVD) have a decreased ability to distinguish between colors, limiting their ability to interact with websites and software packages. User interface designers have taken various approaches to tackle the issue with most offering a high contrast mode. The Web Content Accessibility Guidelines (WCAG) outline some best practices for maintaining accessibility that have been adopted and recommended by several governments; however, it is currently uncertain how this impacts perceived user functionality and if this could result in a reduced aesthetic look. In the absence of subjective data, we aim to investigate how a CVD observer might rate the functionality and aesthetics of existing UIs. However, the design of a comparative study of CVD vs. non-CVD populations is inherently hard, therefore we build on the successful field of physiologically-based CVD models, and propose a novel simulation-based experimental protocol, where non-CVD observers rate the relative aesthetics and functionality of screenshots of 20 popular websites as seen in full color vs. with simulated CVD. Our results show that relative aesthetics and functionality correlate positively and that an operating-system-wide high contrast mode can reduce both aesthetics and functionality. While our results are only valid in the context of simulated CVD screenshots, the approach has the benefit of being easily deployable, and can help to spot a number of common pitfalls in production. Finally, we propose a AAA–A classification of the interfaces we analyzed.
@@ 1 Introduction
Color vision deficiency (CVD, color blindness) is the failure or decreased ability to distinguish between colors under normal illumination. There are over 300 million people with CVD, including approx. 1 in 12 men (8%) and 1 in 250 women (0.5%) birch2012worldwide; rigden1999eye; al2001prevalence. CVD is an X-linked genetic disorder impacting both eyes with varying degrees of prevalence in different populations fakorede2022prevalence. It affects an individual’s ability to perform tasks in both personal and professional settings tagarelli2004colour.
@@
Color is an important asset in user interface (UI) design wcag21, and while the exact impact of color is known to vary between demographics swasty2017does, applications still often rely on established conventions, such as green and red indicating ‘yes‘ and ‘no’ respectively. Objects of the same color satisfy the Gestalt principle of similarity, whereas different colors can help an object stand out or mark figure-ground articulation todorovic2008gestalt. With the ever-increasing color gamut of novel displays wen2021color, new domains are opening up in the use of color; however, some of these domains simply cannot be seen by someone with CVD.
@@
Accessibility is the concept of making UIs equally usable by all types of users, enabling user interactions without barriers. UI designers have the option to support accessibility for CVD users pre-publication or post-publication jefferson2006accommodating, with pre-publication normally resorting to a limited and fixed color palette rigden1999eye, and post-production relying on automatic recoloring iqbal2018adaptive also known as daltonization simon2016evaluating. A hybrid, low-effort approach is to provide support for the operating system’s high contrast mode. The Web Content Accessibility Guidelines (WCAG) wcag21 outline some best practices for accessibility; however, these often only target core functionality. Another consideration for UI is perceived aesthetics. Many designers see aesthetics as inversely proportional to functionality Karvonen_2000, while other evidence points towards a positive correlation between functionality and aesthetics Ahmed_Al; kurosu1995apparent; tractinsky1997aesthetics. In the domain of CVD, a high contrast theme is rarely a top-priority feature, which could imply that people with CVD have a substantially reduced aesthetic experience. However, there is insufficient data to understand CVD users’ perceived functionality and aesthetics of UIs.
A comparative study of UI functionality and aesthetics is inherently challenging. Individuals with CVD cannot judge if a reduced-dimensionality UI they see is equally usable or aesthetic compared to a UI they could never see. Therefore, we instead built on the successful field of physiologically-based CVD simulations and asked 19 non-CVD participants to compare reference UIs to how they might appear to a CVD observer for 20 popular UIs (1449 data points in total). Specifically, we measured mean-opinion scores for functionality and the probability of maintained aesthetics. One clear limitation of our study is non-CVD users’ potential bias towards familiar full-color references, therefore our results cannot be used to conclude whether any of the UIs are fully accessible; however, comparing scores across different applications and accessibility techniques, we can investigate common pitfalls of UI design. Our main contributions can be summarized as follows:
@@
  • Collection of subjective user data from non-CVD observers on the loss of functionality and aesthetics when seen through CVD simulation
@@
The rest of the paper is structured as follows: in Section 2, we discuss the relevant background literature, in Section 3 we describe the proposed experimental methodology, before discussing our findings in Sections LABEL:sec:res and LABEL:sec:discussion.
@@ 2 Background
2.1 CVD
Human color vision is widely considered trichromatic due to the three types of retinal cone cells (L, M, and S) that respond differently to wavelengths of light (Figure 1) wandell1995foundations. According to opponent color theory hurvich1957opponent, higher-order visual functions rely on a differential response of cones, with luminance (L+M) more prominent than chrominance channels (L-M, L+M-S). CVD can occur as a result of cone cells being absent, not working, or having an abnormal response. Severe CVD occurs when two or three types of cone cells are absent (monochromacy, achromatopsia). Less severe CVD occurs when one type of cone is absent (L: protanopia, M: deuteranopia, S: tritanopia), or all three are present, but one cell has an abnormal response (L: protanomaly, M: deuteranomaly, S: tritanomaly) huang2008enhancing. Red-green CVD, specifically protanomaly and deuteranomaly, are the most frequent marey2015ishihara, which can be intuitively explained by the spectral similarity of the M and L cones. This affects approx. 8% and 0.1–0.3% of the male and female population respectively birch2012worldwide.
Color appearance is context-dependent, highly subjective fairchild2013color, and especially as color plays such a key role in visual perception, most find it challenging to imagine how another individual (especially with CVD) might perceive the world around them. However, the physiological mechanisms of early vision are relatively well-understood, and existing models have been shown to be capable of faithfully simulating what a CVD observer might see Machado2009. We build on these models in our experimental methodology.
Figure 1: Normalised spectral sensitivity of the three types of retinal cone cells as a function of light wavelength. Note the similarity of the $M$ and $L$ cones, where minor individual differences can result in an unresolvable overlap, causing observers to have a reduced ability to differentiate between lights in the red-green region. Plotted based on Stockman and Sharpe stockman2000spectral.
@@ 2.2 WCAG
Accessibility is the concept of making UIs equally usable by all types of users, enabling all user interactions without barriers. In our work, we consider functionality to be synonymous with usability, therefore accessibility also aims to maintain functionality. In the past few decades, accessibility has become a legal requirement in several countries easton2012revisiting, with most government agencies recommending the adoption of the World Wide Web Consortium’s (W3C) Web Content Accessibility Guidelines (WCAG) easton2012revisiting; wcag21. In fact, WCAG is consulted by website and UI developers every day to enforce best practices to maintain a level playing field for all users, regardless of age, or physical or mental disabilities. WCAG has also been utilized in academic research to assess website accessibility Al-Khalifa_2010; kuzma2010accessibility; paul2023accessibility.
WCAG 2.1 wcag21, the latest framework, specifies a set of success criteria in an AAAA system, A being the least and AAA being the most restrictive. Minimum accessibility compliance is often set at AA easton2012revisiting. CVD support is addressed across several criteria, e.g.:
@@
It remains uncertain to what extent current UIs implement these recommendations. Also, while the use of contrast ratio thresholds gives designers an intuitive scheme, contrast is defined as a color difference perceived by trichromatic (non-CVD) observers, which is not trivially convertible to contrast for CVD users. Furthermore, WCAG contrast is typically calculated for background vs. foreground, which can fail to capture some impacts of CVD, such as reduced differences across UI components (Figure 2). We accept the importance and impact of WCAG, but we argue that a user study is more appropriate for CVD.
Figure 2: Illustration of a popular quiz platform’s UI processed with simulated tritanomaly Machado2009. While text vs. background contrast remains mostly unchanged, $\Delta E_{2000}$ color difference luo2001development between buttons 1 and 2 is reduced from $25.1$ to $14.1$. Similarly, $\Delta E_{2000}$ between buttons 3 and 4 is reduced from $43.0$ to $18.3$. The colors remain distinguishable, but the difference between them is less noticeable. WCAG analysis would flag an insufficient text vs. background contrast on button 3 (<3), but current guidelines do not address the reduced color difference between buttons.
@@ 2.3 Reducing the impact of CVD
To reduce the impact of CVD, WCAG encourages UI designers to convey information via alternate visual channels. This also agrees with recent research on the use of patterns sajadi2012using, and textures wu2019composition to improve accessibility.
Another widely adopted approach limits the color palette with an increased contrast between key colors rigden1999eye; jefferson2006accommodating. E.g. Windows users can select from multiple operating-system wide high contrast themes that accommodate different types of vision deficiencies. While almost all applications and websites respect the high contrast theme of the operating system, the resulting UI often looks substantially different, with seemingly reduced information for a non-CVD observer (see Figure 3). In our study, we investigate whether this strategy is viable by measuring functionality and aesthetics with the operating system’s high contrast mode enabled and disabled.
An alternative technique is daltonization, an umbrella term for image-based recoloring techniques that enhance displayed images for CVD observers iqbal2018adaptive; ribeiro2019recoloring; machado2010real. Such algorithms aim to alter colors that are isochromatic to CVD observers but are perceived differently by a non-CVD observer. Daltonization could substantially improve CVD accessibility; however, current algorithms are still limited by e.g. run-time cost, lack of temporal coherence, or the ability to accommodate different types and levels of CVD ebelin2023luminance.
2.4 Aesthetics
While accessibility and overall functionality are key features of a good UI, designers, and human-computer interaction (HCI) researchers are increasingly discovering the impact of attractiveness and beauty. An attractive UI can help make a connection between a user and the application turkyilmaz2015user. Interdisciplinary research has investigated the relevance of aesthetics, a classical philosophical field focusing on the “perception of the beautiful in nature and art” oxford_english_dictionary_notitle_2023. A particular focus has been the difference between classical and expressive aesthetics in HCI, with classical aesthetics referring to the traditional notions of an orderly and clear design (e.g. sizing), and expressive aesthetics referring to creativity and originality Hartmann_De. Mahlke and Thüring Mahlke_Thüring_2007 argue that classical aesthetics is perceived more evenly, whereas expressive aesthetics varies with framing; as such, in this paper, we focus on classical aesthetics. For a more complete review of the literature, refer to Ahmed et al. Ahmed_Al.
The relationship between UI aesthetics and functionality has been investigated by a number of authors. Some argue that the two goals often conflict Karvonen_2000, whereas Kurosu and Kashimura kurosu1995apparent found a positive correlation between the beautifulness of ATM interfaces and their usability. Due to the subjective nature of aesthetics, the strength of the correlation coefficient has been shown to be culturally dependent tractinsky1997aesthetics. To our knowledge, no previous publication has investigated aesthetics in the context of CVD. While our simulation-based experimental methodology is not robust enough to measure aesthetics for CVD people, it allows us to hypothesize that functionality and classical aesthetics correlate in CVD-simulated UIs, and that classical aesthetics could be negatively impacted by automatic invasive techniques such as the high contrast mode.
@@ 3 Materials and Methods
In our experiment, we aim to comparatively study the following:
@@
  • Aesthetics: according to classical aesthetics, would the clarity and orderliness of the UI be retained when viewed by a CVD observer?
As discussed earlier, the experimental design is inherently non-trivial. Individuals with CVD cannot make a comparative judgment on whether the stimulus they perceive has the same functionality and aesthetics as a stimulus that they could never see. Conversely, an individual with no CVD does not have sufficient experience to rate how an individual with CVD might perceive a stimulus. A large-scale study with cross-population comparisons of subjective ratings might be considered; however, the subjective nature of aesthetics and the various experiences and consequent biases of the two populations might still prevent such studies from drawing reliable conclusions.
Therefore, we instead build on the successful field of physiologically-based models of CVD simulation that allow non-CVD observers to inspect UIs as seen by CVD observers. Such models are based on a strong scientific understanding of early vision and have been verified by controlled user experiments Machado2009. An inherent limitation of this approach is non-CVD observers’ bias towards the more familiar reference images; therefore, we do not consider the protocol to be suitable to judge whether the functionality and aesthetics would be fully retained for a CVD observer. However, we argue that relative scores across applications and CVD mitigation strategies still reveal crucial insights. Furthermore, a simulation-based pipeline has the benefit that it is easily deployable in a commercial environment.
@@
For an overview of the methodology, see Figure LABEL:fig:methods-fc. In the following subsections, we describe the stimuli, setup, task, participant sample set, and the statistical methods applied.